# The interplay of emission geometry and extragalactic magnetic fields in the spectra of ultra-high-energy cosmic rays

### v1

@@ IV.5 The emission geometry
We also compare the case of an isotropically-emitting source to our benchmark (jetted emission) scenario. This is shown in fig. 12.
@@ V Discussion
[…]
@@
In addition, our results should be interpreted considering the volume filling factor of magnetic fields in-between the sources and Earth. The combined impact of emission geometry and EGMFs on the UHECR spectrum is evidently more pronounced for stronger fields and larger distances. Nevertheless, strong magnetic fields are not volume-filling, such that their influence is limited to only a small subset of propagation paths. Moreover, a detailed three-dimensional model of EGMFs that would allow us to properly model their effect on UHECRs is lacking. [41]. Cosmological magnetohydrodynamics (MHD) simulations and Faraday-rotation measurements consistently indicate that regions with $B\gtrsim 10^{-7}\,\text{G}$ occupy at most a few percent of the comoving volume, being concentrated in galaxy clusters and the densest parts of filaments [42, 43, 44]. As a result, only the spectra of sources embedded in such environments or viewed by an observer embedded therein will be significantly affected [45, 46, 36].
@@
A remarkable consequence of our results is that, even for jets pointing toward us, the spectrum is significantly modified at the highest energies ($E\gtrsim 50\;\text{EeV}$) for coherence lengths comparable to the source distance, as is the case of the upper rows of figures 2, 4, and 8. This effect arises from the extended propagation paths and increased interaction probabilities. This is important for nuclei, since it could affect the mass composition measured at Earth.
@@
The time it takes for CRs to reach Earth from their sources, compared to their distance, also has important implications for multi-messenger observations. For instance, CRs magnetically trapped in galaxy clusters undergo more interactions with the intra-cluster gas when magnetic fields are stronger, consequently increasing the flux of secondary particles such as neutrinos and gamma rays [49, 50, 51]. The same happens in intergalactic space – the longer CRs travel, the more likely they are to produce cosmogenic photons and neutrinos. This could, in principle, completely alter the current estimates of cosmogenic neutrino and photon fluxes [52, 53], significantly increasing them.
Phenomenological fits to the UHECR data [54, 52, 53] are commonly done in one dimension, such that the emission geometry is completely neglected. Nevertheless, even in these simple 1D models, one could think of the trajectory elongation and energy losses as a term that effectively changes the apparent intrinsic properties of the source like $R_{\text{max}}$ and $\alpha$, which then become $R_{\text{max}}=R_{\text{max}}(B,L_{B},D,Z)$ and similarly for the spectral index $\alpha$. In fact, this has been recently done in Ref. [55]. Nevertheless, the hypothesis of isotropic emission implicitly remained. This also has implications for the fluxes of cosmogenic neutrinos and photons discussed mentioned in the previous paragraph, as discussed in Ref. [56].
@@ VI Conclusions
In this work, we investigated the influence of emission geometry and extragalactic magnetic fields on the observed energy spectra of UHECRs, assuming they are emitted in astrophysical jets. Our findings provide novel insights into the conditions under which UHECR events from jets can be observed on Earth, and reveal how the interplay between emission geometry and magnetic fields critically shapes the observed energy spectra.
@@
We have shown that jetted sources emitting towards Earth might not contribute to the UHECR spectrum at specific energy ranges due to a combination of emission geometry, interaction lengths, and magnetic horizons. Off-axis sources whose jets are misaligned with respect to our line of sight can significantly contribute to the observed spectrum, especially at energies $\lesssim 10\;\text{EeV}$. This challenges the intuitive assumption that aligned sources always dominate the flux. It also poses problems for phenomenological models that rely on the hypothesis of isotropic emission.
@@
The data underlying this paper will be shared on reasonable request to the corresponding author.
Acknowledgements
The work of SSS, RMdA and JRTdMN is supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq 406825/2023-8 and Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro - FAPERJ E-26/210.091/2023. RAB is supported by the Agence Nationale de la Recherche (ANR), project ANR-23-CPJ1-0103-01.

### v2

@@ IV.5 Emission geometry
We also compare the case of an isotropically emitting source to our benchmark (jetted emission) scenario. This is shown in Fig. 12.
@@ V Discussion
Some of the scenarios studied here may not be very realistic. For instance, over distances of $\sim 100\;\text{Mpc}$, the contribution of cosmic voids to the volume filling factors of magnetic fields likely increases [36], such that some combinations of parameters such as $R\gtrsim 100\;\text{EeV}$ and $B\gtrsim 10\;\text{nG}$ are not possible in reality, since upper limits for magnetic fields in voids are $47\;\text{pG}$ [21].
@@
In addition, our results should be interpreted considering the volume filling factor of magnetic fields in between the sources and Earth. The combined impact of emission geometry and EGMFs on the UHECR spectrum is evidently more pronounced for stronger fields and larger distances. Nevertheless, strong magnetic fields are not volume filling, such that their influence is limited to only a small subset of propagation paths. Moreover, a detailed three-dimensional model of EGMFs that would allow us to properly model their effect on UHECRs is lacking. [42]. Cosmological magnetohydrodynamics (MHD) simulations and Faraday-rotation measurements consistently indicate that regions with $B\gtrsim 10^{-7}\,\text{G}$ occupy at most a few percent of the comoving volume, being concentrated in galaxy clusters and the densest parts of filaments [43, 44, 45]. As a result, only the spectra of sources embedded in such environments or viewed by an observer embedded therein will be significantly affected [46, 47, 37].
@@
A remarkable consequence of our results is that, even for jets pointing toward us, the spectrum is significantly modified at the highest energies ($E\gtrsim 50\;\text{EeV}$) for coherence lengths comparable to the source distance, as is the case of the upper rows of Figs. 2, 4, and 8. This effect arises from the extended propagation paths and increased interaction probabilities. This is important for nuclei, since it could affect the mass composition measured at Earth.
In our analysis we averaged over several realizations of the turbulent magnetic field. In reality, if only a single realisation were considered, features such as flux magnification or demagnification could appear [50, 51, 52, 53]. Our choice was motivated by computational tractability: sampling the full anisotropic jet emission in combination with multiple EGMF realizations is extremely costly. The adopted approach provides a reasonable first-order approximation to the average UHECR intensity at Earth, but it inevitably suppresses the variance in modification factors that could arise in individual realizations of the field. Such effects are clearly visible in the numerical simulations of Refs. [52, 53], where energy-dependent (de)magnification appears in certain directions.
Causticlike structures and multiple images may be particularly relevant when the source distance is comparable to the coherence length of the magnetic field ($R\sim L_{B}$). In that regime, one expects sharp, direction-dependent amplifications or suppressions of the flux, as well as spectral distortions, which are averaged out here. Nevertheless, this does not compromise the analysis here presented, which demonstrates the initial hypothesis that the UHECR is affected by the interplay of emission geometry and EGMFs. A dedicated study with fixed magnetic-field realizations and true jet emission would be required to capture these lensing effects in detail.
We have also neglected the GMF, which is known to induce causticlike effects [50, 51, 54, 55] depending on the relative strength of the turbulent component with respect to the regular.
@@
The time it takes for CRs to reach Earth from their sources, compared to their distance, also has important implications for multimessenger observations. For instance, CRs magnetically trapped in galaxy clusters undergo more interactions with the intracluster gas when magnetic fields are stronger, consequently increasing the flux of secondary particles such as neutrinos and gamma rays [56, 57, 58]. The same happens in intergalactic space – the longer CRs travel, the more likely they are to produce cosmogenic photons and neutrinos. This could, in principle, completely alter the current estimates of cosmogenic neutrino and photon fluxes [59, 60], significantly increasing them.
Phenomenological fits to the UHECR data [61, 59, 60] are commonly done in one dimension, such that the emission geometry is completely neglected. Nevertheless, even in these simple one-dimensional models, one could think of the trajectory elongation and energy losses as a term that effectively changes the apparent intrinsic properties of the source like $R_{\text{max}}$ and $\alpha$, which then become $R_{\text{max}}=R_{\text{max}}(B,L_{B},D,Z)$ and similarly for the spectral index $\alpha$. In fact, this has been recently done in Ref. [62]. Nevertheless, the hypothesis of isotropic emission implicitly remained. This also has implications for the fluxes of cosmogenic neutrinos and photons discussed mentioned in the previous paragraph, as discussed in Ref. [63].
@@ VI Conclusions
In this work, we investigated the influence of emission geometry and extragalactic magnetic fields on the observed energy spectra of UHECRs, assuming they are emitted in astrophysical jets. Our findings provide novel insights into the conditions under which UHECR events from jets can be observed on Earth and reveal how the interplay between emission geometry and magnetic fields critically shapes the observed energy spectra.
@@
We have shown that jetted sources emitting toward Earth might not contribute to the UHECR spectrum at specific energy ranges due to a combination of emission geometry, interaction lengths, and magnetic horizons. Off-axis sources whose jets are misaligned with respect to our line of sight can significantly contribute to the observed spectrum, especially at energies $\lesssim 10\;\text{EeV}$. This challenges the intuitive assumption that aligned sources always dominate the flux. It also poses problems for phenomenological models that rely on the hypothesis of isotropic emission.
@@
The data that support the findings of this article are not publicly available upon publication beacause it i not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
Acknowledgments
The work of S. S. S., R. M. d. A. and J. R. T. d. M. N. is supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico - CNPq Grant No. 406825/2023-8 and Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro - FAPERJ Grant No. E-26/210.091/2023. R. A. B. is supported by the Agence Nationale de la Recherche (ANR), Project No. ANR-23-CPJ1-0103-01.